# Generative modelling with jump-diffusions

### v1

@@ 6 Conclusion
Extending the noise in the forward process to jump-diffusions leads to a straightforward non-Gaussian generalization of the work of [18]. The results highlight that the score function needs to be adapted to the noise and only takes the conventional form for the Gaussian noise case, while it is otherwise determined by the non-Gaussian details of the noise. Apart from investigating the performance of the framework in more realistic scenarios such as image synthesis, further detailed analysis is also needed to understand the effects of the various parameters and protocols underlying the implementation. One of the shortcomings of the approach is the representation of the generalized score function in terms of an inverse Fourier transform. Instead of using a numerical function approximation, it should also be possible to extend the sliced score matching approach [17] to the jump-diffusion case, which could provide an alternative approach to score estimation.
The fact that the JDL model outperforms LIM and DLPM on such a simple test despite not containing any heavy-tailed characteristics is intriguing. The somewhat poor performance of LIM might be related to the fact that the corresponding SDE neglects an additional noise term due to its finite variation [20]. As such noise and score-function are not exactly balanced as is required by the time reversal formalism and could lead to systematic discrepancies. Further investigation of this interesting theoretical aspect is planned in future work.
@@ Appendix A Time-reversal of jump-diffusion processes
Before we look into more detail into the jump-diffusion generalization of the generative framework outlined in Sec. 2, a heuristic derivation of these results is presented. The aim is to derive an equation for the PDF $\tilde{p}(\mathbf{x},t)$ of the time-reverse or generative process. This PDF can be specified by the condition that $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$ for $t\in[0,T]$. Using the Fokker-Planck equation associated with Eq. (1) we obtain then
@@
Eq. (31) is the Liouville equation corresponding to the deterministic dynamics Eq. (3), which likewise generates the PDF $\tilde{p}(\mathbf{x},t)$ for the initial condition $\mathbf{X}(0)\sim p(\mathbf{x},T)$. So we see that the generative process in the SDE description results from expressing the negative diffusion as a drift plus positive diffusion, while the ODE description results from expressing the negative diffusion as a drift only. We will see that the same heuristic prescription can also be applied for jump diffusions and agrees with rigorous results on the time-reversal of Markov processes with jumps, see Appendix E.
@@
Below, I will derive an explicit expression for $\boldsymbol{\Lambda}$. Eq. (35) is the Liouville equation for the generative process, which takes the form of the probability flow equation
@@
The same result is obtained from rigorous results on the generator for time-reversed Markovian jump processes confirming the validity of Eq. (37), see Appendix E.
@@ A.1 Generalized score function
and likewise in Eq. (39), leading to Eqs. (10,11). The generalized score function is thus identified as $\frac{\nabla\mathcal{V}(\mathbf{x},t)}{p(\mathbf{x},t)}$, which is the central quantity in the jump-diffusion approach to generative modelling.
@@ Appendix B Denoising score matching loss function
In order to derive the jump-diffusion equivalent of the denoising score matching (DSM) loss function Eq. (4), I start with the explicit score matching (ESM) loss function for a given weight function $w(t)$ and follow the derivation of the conventional DSM loss function in [19]:
$\mathcal{J}_{\rm ESM}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\,\mathbb{E}_{\mathbf{Y}(t)}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right\|^{2}\right]$ (46)
@@
$=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\left(\mathbb{E}_{\mathbf{Y}(t)}\left[\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)\|^{2}\right]-2\,\mathbb{E}_{\mathbf{Y}(t)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right]\right)$
$+C_{1}(T)$
since the last term is independent of $\boldsymbol{\theta}$. For the second term I obtain further
@@
where $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$ denotes the FT of the propagator $p(\mathbf{x},t|\mathbf{x}^{\prime})$. Continuing from Eq. (47) yields
@@
$\left.-2\,\mathbb{E}_{\mathbf{Y}(0)}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{F}(\mathbf{Y},t|\mathbf{Y}_{0})}{p(\mathbf{Y},t|\mathbf{Y}_{0})}\right]\right)+C_{2}(T)$
$=$ $J_{\rm ESM}(\boldsymbol{\theta})-C_{1}(T)+C_{2}(T)$
due to Eq. (49) and thus optimizing $\mathcal{J}_{\rm JD}(\boldsymbol{\theta})$ with respect to $\boldsymbol{\theta}$ is equivalent to optimizing $\mathcal{J}_{\rm ESM}(\boldsymbol{\theta})$.
@@ Appendix C The generalized Ornstein-Uhlenbeck process
The forward process is given by the generalized Ornstein-Uhlenbeck process Eq. (15). I first consider an arbitrary jump amplitude distribution $\rho(z)$ with corresponding $\phi(k)$, Eq. (8). The Fokker-Planck equation associated with Eq. (15) is given by setting $\mathbf{f}(\mathbf{x},t)=\frac{1}{2}$ and $g(t)=1$ in Eq. (32), which we want to solve for the conditional PDF $p(\mathbf{x},t|\mathbf{x}^{\prime})$ with the initial condition $p(\mathbf{x},0|\mathbf{x}^{\prime})=\delta(\mathbf{x}-\mathbf{x}^{\prime})$. Fourier-transforming this Fokker-Planck equation yields
@@
$=$ $e^{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}e^{-t/2}}\hat{G}_{1}(k,t)$
@@
$\hat{G}_{1}(k,t)$ $=$ $\exp\left\{-\int_{0}^{t}\mathrm{d}u\,\psi(k\,e^{-u/2})\right\}$ (53)
$=$ $\exp\left\{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)+\lambda\int_{0}^{t}\mathrm{d}u\,\phi(k\,e^{-u/2})\right\}.$ (54)
The propagator is given as the inverse FT of $\hat{p}(\mathbf{k},0|\mathbf{x}^{\prime})$ which can be written in the form
$p(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $G_{1}(|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|,t)$ (55)
@@
$G_{1}(x,t)$ $=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{G}_{1}(k,t).$ (56)
An expression for $\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ can then be directly derived from the definition Eq. (13), for which the FT reads simply
$\hat{\mathcal{V}}(\mathbf{k},t|\mathbf{x}^{\prime})$ $=$ $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})\frac{\psi(k)}{k^{2}}$ (57)
$=$ $e^{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}e^{-t/2}}\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}$
due to Eq. (52). Taking the gradient of the inverse FT yields
$\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $\frac{1}{(2\pi)^{d}}\nabla\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}$ (58)
$=$ $\frac{\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}}{|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|}G_{2}(|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|,t)$ (59)
upon defining the isotropic function
$G_{2}(x,t)=\frac{1}{(2\pi)^{d}}\frac{\partial}{\partial x}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}$ (60)
@@ C.1 Laplace distributed amplitudes
In order to simplify the framework, it is desirable to have a closed analytic form for $\hat{p}(\mathbf{k},0|\mathbf{x}^{\prime})$, which is achieved, for example, by choosing amplitudes according to an isotropic multivariate Laplace distribution $\rho=\text{GL}(0,\sigma^{2}\mathbf{I},1)$, see Appendix D for further details on the multivariate Laplace distribution. The function $\phi(k)$ of Eq. (8) becomes then
@@
$\int_{0}^{t}\mathrm{d}u\,\phi_{\rm L}(k\,e^{-u/2})=\log\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-s}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)$ (62)
Eq. (53) becomes
$\hat{G}_{1}(k,t)=\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}$ (63)
After integrating out the orientational degrees of freedom, the function $G_{1}(x,t)$, Eq. (56), can be expressed as the one-dimensional integral
$G_{1}(x,t)$ $=$ $\frac{1}{(2\pi)^{d/2}}\int_{0}^{\infty}\mathrm{d}k\,k\,\left(\frac{k}{x}\right)^{\frac{d}{2}-1}J_{\frac{d}{2}-1}(k\,x)\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}$ (64)
where $J_{\mu}$ denotes the Bessel function of the first kind. For $t\to\infty$ the stationary distribution $p_{\rm st}(x)=\lim_{t\to\infty}p(\mathbf{x},t|\mathbf{x}_{0})$ is obtained as
$p_{\rm st}(x)$ $=$ $\lim_{t\to\infty}G_{1}(x,t)$ (65)
$=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\left(\frac{1}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}e^{-\frac{D}{2}k^{2}}$
The function $G_{2}(x,t)$, Eq. (60), can likewise be expressed as a one-dimensional integral:
$G_{2}(x,t)$ $=$ $\frac{1}{(2\pi)^{d/2}}\frac{\partial}{\partial x}\int_{0}^{\infty}\mathrm{d}k\,k\,\left(\frac{k}{x}\right)^{\frac{d}{2}-1}J_{\frac{d}{2}-1}(k\,x)\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}$ (66)
$=$ $-\frac{x}{(2\pi)^{d/2}}\int_{0}^{\infty}\mathrm{d}k\,k\,\left(\frac{k}{x}\right)^{\frac{d}{2}}J_{\frac{d}{2}}(k\,x)\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}\times$
$e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}\left(\frac{D}{2}-\lambda\frac{1}{k^{2}\left(1+\frac{\sigma^{2}}{2}k^{2}\right)}+1\right)$
substituting $\psi(k)$ of Eq. (7) with Eq. (61).
C.2 The conditional generalized score function for $\psi(k)=k^{\alpha}$
With Eqs. (58,53) one obtains for the special case $\psi(k)=k^{\alpha}$
@@ C.2 The conditional generalized score function for $\psi(k)=k^{\alpha}$
$\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $\frac{1}{(2\pi)^{d}}\nabla\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}e^{-k^{\alpha}f(t)}k^{\alpha-2}$ (67)
$=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}(-i\mathbf{k})e^{-k^{\alpha}f(t)}k^{\alpha-2}$
$=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}i\nabla_{\mathbf{k}}\frac{e^{-k^{\alpha}f(t)}}{\alpha\,f(t)}$
$=$ $-\frac{\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}{\alpha\,f(t)}\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}e^{-k^{\alpha}f(t)}$
@@
$=$ $-\frac{\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}{\alpha\,f(t)}p(\mathbf{x},t|\mathbf{x}^{\prime})$
where $f(t)=\int_{0}^{t}\mathrm{d}u\,e^{-\alpha\,u/2}=\frac{2}{\alpha}\left(1-e^{-t}\right)$. Thus the conditional generalized score function becomes Eq. (27).
Appendix D The multivariate generalized Laplace distribution

### v2

@@ 6 Conclusion
Extending the noise in the forward process to jump-diffusions leads to a straightforward non-Gaussian generalization of the standard DM framework of [24], which is likewise based on a MSE loss. The results highlight that the score function needs to be adapted to the noise and only takes the conventional form for the Gaussian noise case, while it is otherwise determined by the non-Gaussian details of the noise. The implementation of the JL model rivals both the LIM and standard Gaussian model in its analytical and computational simplicity and might lead to performance improvements in specific applications. In particular the fact that it performs quite well for a small number of discretization steps as shown in Fig. 1(c,f) seems promising. Apart from investigating the performance of the framework in more realistic scenarios such as image synthesis, further detailed analysis is also needed to understand the effects of the various parameters and protocols underlying the implementation. It should be highlighted that the design space of optimal protocols and network architectures is likely different from that of Gaussian models, so a lot more work is needed to realize the potential impact of non-Gaussian models.
On the theoretical side, in addition to extending the framework to more general Markov processes conceptually, it would be interesting to investigate other specific jump amplitude distributions and the combined effect of Gaussian white noise and jumps in detailed implementations. While the generalized Ornstein-Uhlenbeck process is generally integrable in Fourier space (see Appendix C) the resulting expressions for the conditional PDF and generalized score function are typically not available in closed form. It might thus be worthwhile to consider other score matching approaches such as, e.g., sliced score matching [23] and investigate how these can be adapted to the non-Gaussian case.
@@ Appendix A Time-reversal of jump-diffusion processes
Before we look in more detail into the jump-diffusion generalization of the standard generative framework with Gaussian noise outlined in Sec. 2, a heuristic derivation of these results is presented. The aim is to derive an equation for the PDF $\tilde{p}(\mathbf{x},t)$ of the time-reverse or generative process. This PDF can be specified by the condition that $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$ for $t\in[0,T]$. Using the Fokker-Planck equation associated with Eq. (1) we obtain then
@@
Eq. (40) is the continuity equation corresponding to the deterministic dynamics Eq. (3), which likewise generates the PDF $\tilde{p}(\mathbf{x},t)$ for the initial condition $\mathbf{X}(0)\sim p(\mathbf{x},T)$. So we see that the generative process in the SDE description results from expressing the negative diffusion as a drift plus positive diffusion, while the ODE description results from expressing the negative diffusion as a drift only. We will see that the same heuristic prescription can also be applied for jump diffusions and agrees with rigorous results on the time-reversal of Markov processes with jumps, see Appendix G.
@@
Below in Eq. (A.1), I will derive an explicit expression for $\boldsymbol{\Lambda}$. Eq. (44) is the continuity equation for the generative process, which takes the form of the probability flow ODE
@@
The same result is obtained from rigorous results on the generator for time-reversed Markovian jump processes confirming the validity of Eq. (46), see Appendix G.
@@ A.1 Generalized score function
Identifying the generalized score function as in Eq. (9) thus leads to Eq. (11). Likewise, Eq. (48) leads to Eq. (12).
@@ Appendix B Denoising score matching loss function
In order to derive the jump-diffusion equivalent of the denoising score matching (DSM) loss function Eq. (4), I start with the explicit score matching (ESM) loss function for the generalized score function $\mathcal{S}(\mathbf{x},t)$, Eq. (9), and follow the derivation of the conventional DSM loss function in [26]:
$\mathcal{J}_{\rm ESM}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\,\mathbb{E}_{\mathbf{Y}(t)}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\mathcal{S}(\mathbf{x},t)\right\|^{2}\right]$
$=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\,\mathbb{E}_{\mathbf{Y}(t)}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right\|^{2}\right]$
@@
$=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\left(\mathbb{E}_{\mathbf{Y}(t)}\left[\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)\|^{2}\right]-2\,\mathbb{E}_{\mathbf{Y}(t)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right]\right)+C_{1}(T)$
since the last term is independent of $\boldsymbol{\theta}$. The function $w(t)$ is a given weight function. For the second term I obtain further
@@
where $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$ denotes the FT of the conditional PDF $p(\mathbf{x},t|\mathbf{x}^{\prime})$. Continuing from Eq. (56) yields
@@
$\left.-2\,\mathbb{E}_{\mathbf{Y}(0)}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{Y},t|\mathbf{Y}_{0})}{p(\mathbf{Y},t|\mathbf{Y}_{0})}\right]\right)+C_{2}(T)$
$=$ $\mathcal{J}_{\rm ESM}(\boldsymbol{\theta})-C_{1}(T)+C_{2}(T)$
due to Eqs. (LABEL:jesm,58) and thus optimizing $\mathcal{J}_{\rm JD}(\boldsymbol{\theta})$ with respect to $\boldsymbol{\theta}$ is equivalent to optimizing $\mathcal{J}_{\rm ESM}(\boldsymbol{\theta})$.
@@ Appendix C The generalized Ornstein-Uhlenbeck process
The forward process under consideration in the detailed implementation is a generalized Ornstein-Uhlenbeck process defined by Eqs. (1,6) with
$\mathbf{f}(\mathbf{x},t)=-\frac{1}{2}\,\mathbf{x}\quad,\qquad\qquad g(t)=1$ (60)
I first consider an arbitrary jump amplitude distribution $\rho$ with corresponding $\phi(k)$, Eq. (8). The Fokker-Planck equation associated with Eqs. (1,6) is given in Eq. (41), which we want to solve for the conditional PDF $p(\mathbf{x},t|\mathbf{x}^{\prime})$ with the initial condition $p(\mathbf{x},0|\mathbf{x}^{\prime})=\delta(\mathbf{x}-\mathbf{x}^{\prime})$. Fourier-transforming this Fokker-Planck equation for the choice of $\mathbf{f},g$ in Eq. (60) yields
@@
$=$ $e^{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}e^{-t/2}}\hat{G}(k,t)$
@@
$\hat{G}(k,t)$ $=$ $\exp\left\{-\int_{0}^{t}\mathrm{d}u\,\psi(k\,e^{-u/2})\right\}$ (63)
$=$ $\exp\left\{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)+\lambda\int_{0}^{t}\mathrm{d}u\,\phi(k\,e^{-u/2})\right\}$ (64)
The conditional PDF is given as the inverse FT of $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$ which can be written in the form
$p(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $G(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2},t)$ (65)
@@
$G(\mathbf{x},t)$ $=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{G}(k,t).$ (66)
The definition Eq.(14) implies that
$\mathcal{F}\left\{\frac{\partial}{\partial x_{i}}\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})\right\}=-i\,k_{i}\,e^{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}e^{-t/2}}\hat{G}(k,t)\frac{\psi(k)}{k^{2}}$ (67)
Moreover, Eq. (63) leads to the exact relationship
$\frac{\partial}{\partial k_{i}}\hat{G}(k,t)=2\frac{k_{i}}{k^{2}}\left(\psi(k\,e^{-t/2})-\psi(k)\right)\hat{G}(k,t)$ (68)
and thus
$k_{i}\,\hat{G}(k,t)\frac{\psi(k)}{k^{2}}=-\frac{1}{2}\frac{\partial}{\partial k_{i}}\hat{G}(k,t)+k_{i}\,\hat{G}(k,t)\frac{\psi(k\,e^{-t/2})}{k^{2}}$ (69)
While not much has been gained here, the last equation simplifies further for Laplace distributed amplitudes.
@@ C.1 Laplace distributed amplitudes
In the following I set $\lambda=1$ and focus on jumps according to an isotropic multivariate Laplace distribution $\text{L}_{d}(\sigma^{2})$, which has the characteristic function Eq. (87) (see Appendix D). The function $\phi$ of Eq. (8) becomes then
@@
$\int_{0}^{t}\mathrm{d}u\,\phi_{\rm L}(k\,e^{-u/2})=\log\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)$ (71)
As a consequence, $\hat{G}(k,t)$ of Eq. (63) assumes the form
$\hat{G}(k,t)$ $=$ $\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}$ (72)
$=$ $\left(\frac{1-e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}+e^{-t}\right)e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}$ (73)
Since $\psi(k)=\frac{D}{2}k^{2}-\phi_{\rm L}(k)$, we also obtain
$\frac{\psi(k\,e^{-t/2})}{k^{2}}=\frac{D}{2}e^{-t}+\frac{\frac{\sigma^{2}}{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}$ (74)
Multiplying Eq. (74) with Eq. (72), the last term in Eq. (69) now becomes
$\hat{G}(k,t)\frac{\psi(k\,e^{-t/2})}{k^{2}}$ $=$ $\frac{D}{2}e^{-t}\hat{G}(k,t)+\frac{\frac{\sigma^{2}}{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}$ (75)
$=$ $\frac{D}{2}e^{-t}\hat{G}(k,t)+\frac{\frac{\sigma^{2}}{2}e^{-t}}{1-e^{-t}}\left(\hat{G}(k,t)-e^{-t}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}\right)$ (76)
where in the last step Eq. (73) has been used. Substituting Eq. (76) in Eq. (69) and subsequently Eq. (67) yields
@@
$\mathcal{F}\left\{\frac{\partial}{\partial x_{i}}\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})\right\}$ $=$ $i\,e^{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}e^{-t/2}}\left(\frac{1}{2}\frac{\partial}{\partial k_{i}}\hat{G}(k,t)-k_{i}\left(\frac{D}{2}e^{-t}+\frac{\frac{\sigma^{2}}{2}e^{-t}}{1-e^{-t}}\right)\hat{G}(k,t)\right$. (77)
$\left.+\,k_{i}\,\frac{\frac{\sigma^{2}}{2}e^{-2t}}{1-e^{-t}}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}\right)$
The inverse FT is now exact
@@
$\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $-\frac{1}{2}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)p(\mathbf{x},t|\mathbf{x}^{\prime})+\left(\frac{D}{2}e^{-t}+\frac{\frac{\sigma^{2}}{2}e^{-t}}{1-e^{-t}}\right)\nabla p(\mathbf{x},t|\mathbf{x}^{\prime})$ (78)
$-\frac{\frac{\sigma^{2}}{2}e^{-2t}}{1-e^{-t}}\nabla\mathcal{N}(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2},D(1-e^{-t})\mathbf{I}_{d})$
where $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ denotes the $d$-dimensional multivariate normal distribution with mean vector $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Unfortunately, $p(\mathbf{x},t|\mathbf{x}^{\prime})$ as inverse FT of Eq. (73) can not be derived in closed analytical form. In order to have an easily applicable generative model, I further consider the pure jump case $D=0$.
C.2 The JL model
Setting $D=0$ yields in Eq. (72)
$\hat{G}(k,t)$ $=$ $\frac{1-e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}+e^{-t}$ (79)
which has the exact inverse FT (see Appendix D)
$G(\mathbf{x},t)$ $=$ $\rho_{\rm L}(\mathbf{x},\sigma)\left(1-e^{-t}\right)+\delta(\mathbf{x})e^{-t}$ (80)
where $\rho_{\rm L}(\mathbf{x},\sigma)$ is the distribution $\text{L}_{d}(\sigma^{2})$ itself, see Eq. (89). The conditional PDF is thus determined with Eq. (65).
In addition, setting $D=0$ in Eq. (78) yields
$\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $-\frac{1}{2}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)p(\mathbf{x},t|\mathbf{x}^{\prime})$ (81)
$+\left(\frac{\frac{\sigma^{2}}{2}e^{-t}}{1-e^{-t}}\right)\nabla\left(p(\mathbf{x},t|\mathbf{x}^{\prime})-\delta\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)e^{-t}\right)$
$=$ $-\frac{1}{2}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)p(\mathbf{x},t|\mathbf{x}^{\prime})+\frac{\sigma^{2}}{2}e^{-t}\nabla\rho_{\rm L}(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2},\sigma)$
$=$ $-\frac{1}{2}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)p(\mathbf{x},t|\mathbf{x}^{\prime})$
$+\frac{\sigma^{2-d}\,e^{-t}}{(2\pi)^{d/2}}\nabla\left(\frac{\|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\|}{\sqrt{2}\,\sigma}\right)^{1-\frac{d}{2}}K_{1-\frac{d}{2}}\left(\frac{\|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\|}{\sigma/\sqrt{2}}\right)$
After evaluating the derivatives of the Bessel function and simplifications, we obtain the conditional generalized score function in the form
$\mathcal{S}(\mathbf{x},t)=\frac{\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})}{p(\mathbf{x},t|\mathbf{x}^{\prime})}=\frac{\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}}{\|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\|}\mathcal{G}(\|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\|,t)$ (82)
where $\mathcal{G}(x,t)$ is defined in Eq. (20).
C.3 Variance of the score estimator
Figure 2: The lhs of Eq. 83 obtained by Monte-Carlo sampling and divided by $d$. The increments $\mathbf{J}(t)$ are sampled as outlined in Sec. 3.3.
In the denoising score matching loss function Eq. (25) one can identify the score estimator $\frac{\mathbf{J}(t)}{\|\mathbf{J}(t)\|}\hat{\mathcal{G}}(\|\mathbf{J}(t)\|,t)$, where $\hat{\mathcal{G}}(x,t)=(1-e^{-t})\mathcal{G}(x,t)$ and the jump increments $\mathbf{J}(t)$ have the distribution Eq. (30). While the variance of this estimator can not be evaluated analytically, it is straightforward to compute it by sampling. As Fig. 2 shows, for constant $t$ the variance
$\sum_{j=1}^{d}{\rm Var}\left(\frac{J_{i}(t)}{\|\mathbf{J}(t)\|}\hat{\mathcal{G}}(\|\mathbf{J}(t)\|,t))\right)\propto d$ (83)
then scales linearly in $d$ just like in the Gaussian case. In this case, the estimator is just the noise itself $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, see Eq. 96.
Appendix D The multivariate Laplace distribution